I am 9th std student preparing for IIT-JEE.


Please send me the proof of heron's formula

Grade:9

1 Answers

T.S. Naveen
13 years ago

A Proof of the Pythagorean Theorem From Heron's Formula

Let the sides of a triangle have lengths a,b and c. Introduce the semiperimeter p = (a + b + c)/2 and the area S. Then Heron's formula asserts that

  S2 = p(p - a)(p - b)(p - c)

W. Dunham analyzes the original Heron's proof in his Journey through Genius.

For the right triangle with hypotenuse c, we have S = ab/2. We'll modify the right hand side of the formula by noting that

  p - a = (- a + b + c)/2, p - b = (a - b + c)/2, p - c = (a + b - c)/2

It takes a little algebra to show that

 
16S2 = (a + b + c)(- a + b + c)(a - b + c)(a + b - c)
  = 2a2b2 + 2a2c2 + 2b2c2 - (a4 + b4 + c4)

For the right triangle, 16S2 = 4a2b2. So we have

  4a2b2= 2a2b2 + 2a2c2 + 2b2c2 - (a4 + b4 + c4)

Taking all terms to the left side and grouping them yields

  (a4 + 2a2b2 + b4) - 2a2c2 - 2b2c2 + c4 = 0

With a little more effort

  (a2 + b2)2 - 2c2(a2 + b2) + c4 = 0

And finally

  [(a2 + b2) - c2]2 = 0

Remark

For a quadrilateral with sides a, b, c and d inscribed in a circle there exists a generalization of Heron's formula discovered by Brahmagupta. In this case, the semiperimeter is defined as p = (a + b + c + d)/2. Then the following formula holds

  S2 = (p - a)(p - b)(p - c)(p - d)

Since any triangle is inscribable in a circle, we may let one side, say d, shrink to 0. This leads to Heron's formula.